OpenStudy (anonymous):

A circle has a circumference of 48π meters. Find the radius, r, of the circle. (C = 2πr)

OpenStudy (anonymous):

\(C=2 \pi r\) rearrange this equation in terms of the radius you'll have \(r=\frac{C}{2 \pi}\) substitute C for 48π, and you'll get the answer. welcome to openstudy!:)

OpenStudy (anonymous):

so just write 48/2 pi?

OpenStudy (anonymous):